an electrician can install the electric wires in a house in 14 hours.a second electrician requires 18 hours. how long would it take both electricians, working together, to install the wires?
Answer by edjones(7569)

1st does 1/14 of the job in an hour
2nd does 1/18 of the job in an hour
1/14+1/18
least common multiple=126: 9/126+7/126=16/126 part of the job done in 1 hr by both of them.
if we flip the 16/126 over we get how long the job will take: 126/16= 7 7/8 hrs.
Ed